Find the general solution for dy/dx = (x+y+5)^2
Answer #1

Lit z = x+y+5. then e de So do y = (x+y+5)? | 7 dienos - 1 = za - dx. Integreting, tantz=xte (e-constant) za tan (x+C) - X++